OpenStudy (anonymous):

A piece of wire 40cm long is cut into two pieces. One piece is bent into the shape of a square and the other into the shape of a circle. How should the wire be cut so that the total enclosed is (a) maximum? (b)minimum?

OpenStudy (anonymous):

let 40 = s + c A_square = (s/4)^2 A_circle = pi(c/(2pi))^2 to make a function of 1 variable solve 40 = s + c for either s or c and substitute in the proper area equation.

OpenStudy (anonymous):

what do u think it is

OpenStudy (anonymous):

you can then max and min using calculus (derivative)

OpenStudy (anonymous):

So if I isolate the c, which would be 40-s=c, I would sub it into the area_circle?

OpenStudy (anonymous):

yep then the area is the sum of the areas (square and circle)

OpenStudy (anonymous):

find dA/ds and set equal to 0. hopefully you have 2 zeros, one which will max and one which will min. but you will have to check

OpenStudy (anonymous):

did you figure it out?

OpenStudy (anonymous):

I figured out the circle one, which is 40cm.

OpenStudy (anonymous):

I didn't quite get the square one...

OpenStudy (anonymous):

A = A_square + A_circle you have to maximize and minimize A not the individual area functions

OpenStudy (anonymous):

@pgpilot326 sorry, I didn't quite it...

OpenStudy (anonymous):

get*

OpenStudy (anonymous):

A = (s/4)^2 + pi((40-s)/2pi)^2 find dA/ds and then find the zeros of dA/ds check each zero to see if it maximizes or minimizes the function then you can decide how to cut, accordingly

OpenStudy (anonymous):

are you good now or do you have more questions on this?

OpenStudy (anonymous):

Yes, I think I got it now. Thanks!

OpenStudy (anonymous):

right on. good luck!